Some systems are under maintenance.

ETH-Services Status

Status of AS200462

  • Network

    Last updated 2025-09-26 20:44:59

    Operational
  • Power

    Last updated 2025-09-18 17:23:01

    Operational
  • Cooling

    Last updated 2025-09-18 17:23:01

    Operational
  • AS174 - Cogent

    Last updated 2025-12-21 18:01:08

    Operational
  • AS1299 - Arelion

    Last updated 2026-01-31 21:27:50

    Operational
  • AS2914 - NTT

    Last updated 2026-01-20 03:34:47

    Operational
  • AS5405 - Inter.Link

    Last updated 2026-02-03 14:40:19

    Under maintenance
  • AS5511 - Orange

    Last updated 2026-01-26 13:21:58

    Operational
  • AS6453 - TATA

    Last updated 2025-12-21 18:01:09

    Operational
  • AS6762 - Telecom Italia Sparkle

    Last updated 2026-01-30 21:22:40

    Operational
  • AS6830 - Liberty Global

    Last updated 2025-11-11 22:18:39

    Operational
  • AS9002 - RETN

    Last updated 2025-12-12 15:06:02

    Under maintenance
  • AS34549 - Meerfarbig

    Last updated 2025-11-11 22:18:39

    Operational
  • Cloud

    Last updated 2025-09-18 17:23:01

    Operational
  • DNS

    Last updated 2025-10-23 12:38:37

    Operational
  • DDoS-Filter

    Last updated 2025-09-18 17:23:01

    Operational
  • E-Mail

    Last updated 2025-09-18 17:23:01

    Operational
  • VServer

    Last updated 2025-09-18 17:23:01

    Operational
  • VServer - Backups

    Last updated 2025-09-18 17:23:01

    Operational
  • Web

    Last updated 2025-09-18 17:23:01

    Operational

Past Incidents

Fixed

2 weeks ago —

Sparkle has implemented various changes that resolve the situation. BGP has been enabled again, and the situation is stable. We will monitor the situation carefully and intervene when necessary.

Investigating

2 weeks ago —

The latest update remains in place. Sparkle NOC applied various changes, which remain unsuccessful. The carrier remains drained until restoration is in place. No customer impact is expected at this point.

Reported

2 weeks ago —

After further communication, Sparkle installed another faulty configuration change, resulting in blackholing of egress traffic. Currently, BGP is disabled and traffic is flowing via alternative paths to mitigate the impact.

Customers traversing the Sparkle backbone may have experienced packet loss between 10:30 UTC and 12:30 UTC.

We are in constant contact with the Sparkle NOC to address the issue, and we apologize for the inconvenience caused.

Fixed

2 weeks ago —

Sparkle has reverted the faulty changes and sessions have been enabled again.

2 weeks ago —

We are currently observing a partial null route of traffic via Carrier Sparkle due to a false configuration change on Sparkle’s side. BGP has been drained, and the impact is mitigated for the moment. We will get in touch with the Sparkle NOC for a permanent solution.